# A short trajectory is all you need: A transformer-based model for long-time dissipative quantum dynamics

### v1

@@ Abstract
Data sets for training, validation, and testing. The data set used here is taken from Ref. 31. We describe it here for completeness. The reduced density matrix of the spin-boson Hamiltonian given by Eq. (1) is calculated for all combinations of the following parameters: $\epsilon/\Delta=\{0,1\}$, $\lambda/\Delta=\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0\},\;\omega_{c}/\Delta=\{1,2,3,4,5,6,7,8,9,10\}$, and $\beta\Delta=\{0.1,0.25,0.5,0.75,1\}$. HEOM method implemented in QuTiP software packageJohansson, Nation, and Nori (2012) was used in all calculations. The total propagation time was $t_{\max}\Delta=20$. The integration time-step was set to $t\Delta=0.05$. In total, 1000 HEOM calculations, 500 for symmetric $(\epsilon/\Delta=0)$ and 500 for asymmetric $(\epsilon/\Delta=1)$ spin-boson Hamiltonian were performed. Time-evolved subsystem reduced density matrices were saved every $\mathrm{d}t\Delta=0.1$. The population difference $\left\langle\hat{\sigma}_{z}(t)\right\rangle$ are calculated from the reduced denisty matrices and processed into shorter sequences of length $T$ by window slicing. Herrera Rodríguez and Kananenka (2021); Ullah and Dral (2021); Lin et al. (2021) Namely, for a time series $\mathbf{x}=\left(x^{(1)},\ldots,x^{(L)}\right)$, where $\left\langle\hat{\sigma}_{z}(t^{(j)})\right\rangle$ is denoted by $x^{(j)}$ for compactness, a slice is a subset of the original time series defined as $\mathbf{s}_{i:j}=\left(x^{(i)},\ldots,x^{(j)}\right)$ where $1\leqslant i\leqslant j\leqslant P$. For a given time series $\mathbf{x}$ of length $L$, and the length of the slice $P$, a set of $L-P+1$ sliced time series $\left\{\mathbf{s}_{1:P},\mathbf{s}_{2:P+1},\ldots,\mathbf{s}_{L-P+1:L}\right\}$ was generated. This procedure was also applied to the times of the sequence $\mathbf{t}$. Finally, the total data set $\mathcal{D}=\left\{\left(\{\mathbf{x}_{i},\mathbf{t}_{i}\},y_{i}\right)\right\}_{i=1}^{N}$ containing time series $\mathbf{x}_{i}$ and $\mathbf{t}_{i}$ with their corresponding labels $y_{i}$ is obtained by setting $1,\ldots,T$ elements of each slice, with $T=P-1$, to an input time-series $\mathbf{x}_{i}$ and the last ($P$ th) element of each slice to the associated label $y_{i}$.
@@
In general, the size of the window $P-1$, or equivalently $T$, should be treated as a hyperparameter but, following our previous works, Herrera Rodríguez and Kananenka (2021); Rodríguez et al. (2022) we set $T=0.2L$. Window slicing is applied to all 1000 HEOM reduced density matrices with different system and system-bath parameters. For each set of parameters, the initially calculated set of time-evolved $\left\langle\hat{\sigma}_{z}(t)\right\rangle$ with $L=t_{\max}/\mathrm{d}t=200$ generates 160 data points for the data set with $T=41$ (including $t\Delta=0$ point).
@@
From the raw HEOM data set of 1000 trajectories, 100 randomly chosen trajectories are taken as the hold-out test set, which is used for testing and generating the results. The remaining set of 900 trajectories is transformed into 144000 trajectories by window slicing described above. In total 72000 short-time $\left\langle\hat{\sigma}_{z}(t)\right\rangle$ trajectories for symmetric and 72000 for asymmetric spin-boson models were generated. Each trajectory has a time length of $t\Delta+\mathrm{d}t\Delta=4.1$ where $t\Delta=4.0$ is a part of a trajectory used as input vector $\mathbf{x}_{i}$ and the last point $t\Delta+\mathrm{d}t\Delta=4.1$ is used as a label $y_{i}$. Following previous similar works the input data was not normalized.Herrera Rodríguez and Kananenka (2021); Ullah and Dral (2021) This data set of supervised trajectories is the training set which is randomly partitioned into two subsets: a sub-training set, which contains $80\%$ of the data and a validation set containing $20\%$ of the data. Transformer model parameters, weights and biases, were initially fitted on the sub-training set and the validation set is used for monitoring the performance of the models (mainly to prevent overfitting).
@@
Transformer Model. The model architecture used in this work is shown in Fig. 3. The input is the fixed-length trajectories corresponding to the population difference x and the corresponding time values t. In order to use the self-attention mechanism, the inputs are embedded in objects $X\in\mathbb{R}^{T\times d_{p}}$ that represent the original information of the time series in a $d_{p}$ dimensional space. After the self-attention block the input $X$ will be transformed in $Z\in\mathbb{R}^{T\times d_{p}}$. The population difference elements are projected into the $d_{p}=1\times 64$ space using a linear point-wise feed-forward neural network (PW-FFNN) without hidden layers. This operation transforms the population difference into a vector of size $d_{p}$. The time values are projected using the PE. At each step the PE is calculated using Eq. (8) projecting the time value into a vector of $d_{p}$. Note the match between the projected dimension of the FFNN for the population difference and the PE, doing so the two inputs are encoded in the space of the same dimensionality, and can be added as a single input for the self-attention layer. The FFNN should be able to learn how to encode the values of the population difference without interfering with the constant part of the PE. This means that the population difference will be represented in the higher dimension of $X=P+PE$ since the higher dimensions of the PE are constant as shown in Fig 2. After the addition, the inputs will be represented by the matrix $X$ of $41\times 64$.
@@
After the transformer layers, in order to reduce the dimensions before flattening, a PW-FNN is applied. After the flattening, two fully connected FFNNs are used with a number of neurons of 1024 and 1408 respectively and RELU as the activation function. The final output of the ML model is a single neuron that gives the population difference at the next time $\hat{\sigma}(T+1)$.
@@
The model was implemented using Keras Chollet et al. (2015) software with TensorFlow blackened.Abadi et al. (2015) The model was trained using the adaptive moment estimation (Adam) algorithm Diederik (2014) with an initial learning rate of $1.0\times 10^{-4}$. The batch size was set to $N_{b}=128$ and the mean square error (MSE) was used as the loss function. The hyperparameters of the model: ##\##heads, $d_{p}$, number transformer layers, and neurons in the fully-connected layers were optimized on the hold-out set for the asymmetric spin-boson model using KerasTunerO’Malley et al. (2019) for 50 iterations of Bayesian optimization. 500 epochs were used in the training. After the best model was chosen it was further trained for up to 4500 epochs and the model with the lowest validation loss during the training was used. The MSE of the validation set was below $10^{-6}-10^{-7}$. This model was used to predict the population dynamics for the asymmetric spin-boson model presented below.
@@
To generate the population dynamics for the symmetric spin-boson model, we retrained the model described above (for the asymmetric spin-boson model) based on the data set for the symmetric spin-boson Hamiltonian. The same hyperparameters and the initial weights were taken from the already trained asymmetric model. Since the asymmetric model already learned a good representation of the input trajectory for the asymmetric model, the symmetric ML model was trained only for 1000 epochs and the best model was chosen again based on the validation MSE. The training was done on an NVIDIA RTX A6000 Graphic Card.
Figure 4: Expectation values $\left\langle\hat{\sigma}_{z}(t)\right\rangle$ for asymmetric spin-boson model with $\epsilon=1$ as a function of time. Results predicted by the transformer-based ML model developed in this work (red circles) are compared to the numerically exact HEOM results (blue) for the following parameters: (a) $\lambda=0.1,\omega_{c}=6.0,\beta=0.75$; (b) $\lambda=0.3,\omega_{c}=8.0,\beta=1.0$; (c) $\lambda=0.2,\omega_{c}=10.0,\beta=0.25$; (d) $\lambda=0.4,\omega_{c}=8.0,\beta=0.75$; (e) $\lambda=0.8,\omega_{c}=10.0,\beta=1.0$; (f) $\lambda=0.7,\omega_{c}=10.0,\beta=0.1$. All parameters are in the units of $\Delta$.
@@
Starting from the input sequence of $\mathbf{x}=\left(x^{(1)},\ldots,x^{(T)}\right)$ one uses the ML model to generate the population difference at $T+1$ (or $P$th) time step $y$. The newly predicted population difference is then combined with $T$ previous population differences to form another input vector $\mathbf{x}=\left(x^{(2)},\ldots,x^{(T)},y\to x^{(T+1)}\right)$ which is fed into the model again to generate the next prediction $y\to x^{(T+2)}$. Repeating this procedure will generate the long-time population dynamics from the initial short-time input. It should be noted that the next time value should match the time sequence, i.e. any time point should be $t^{(n)}=(n-1)*dt$, and it is necessary to update the time input correspondingly.
In Fig. 4 we show the predictions of the population difference of our transformer model for the asymmetric spin-boson system for six representative sets of parameters. The agreement between ML-generated dynamics and numerically exact HEOM dynamics is excellent. The average error (MAE) between the predicted and reference trajectories for the entire hold-out set of 50 diverse reduced density matrix trajectories was $\epsilon=7.45\times 10^{-3}$. We stress that only a short trajectory of length $t\Delta=4$ is used as an input and these trajectories were not in the training data. To put this value into perspective, the error achieved in this work with the transformer architecture is lower than the lowest error for any neural-network-based models trained, for the same system, in our previous comprehensive study of ML models for long-time quantum dissipative dynamics.Rodríguez et al. (2022) There the lowest error of $\epsilon=2.14\times 10^{-2}$ was achieved for a convolutional gated recurrent unit (GRU) model. In that work we also established that KRR models can be more accurate than their neural network counterparts. Importantly, the performance of our transformer-based model is now at the level of KRR models yet by using neural networks we can avoid the known issues associated with KRR models such as poor scaling with the amount of the training data. However, it is important to note that the number of trainable parameters in the case the transformer-based model presented here is $1,918,018$ which is four times the number of parameters of the convolutional GRU model. Increasing the number of parameters, however, does not necessarily guarantee the increase in the performance of the model.
@@
Figure 5: Expectation values $\left\langle\hat{\sigma}_{z}(t)\right\rangle$ for symmetric spin-boson model as a function of time. Results predicted by the transformer-based ML model developed in this work (red circles) are compared to the numerically exact HEOM results (blue) for the following parameters: (a) $\lambda=0.2,\omega_{c}=8.0,\beta=1.0$; (b) $\lambda=0.4,\omega_{c}=10.0,\beta=1.0$; (c) $\lambda=0.2,\omega_{c}=10.0,\beta=0.25$; (d) $\lambda=0.1,\omega_{c}=4.0,\beta=0.1$; (e) $\lambda=0.8,\omega_{c}=3.0,\beta=1.0$; (f) $\lambda=1.0,\omega_{c}=2.0,\beta=0.1$. All parameters are in the units of $\Delta$.
@@
Fig. 5 illustrates the results for the symmetric spin-boson model. Similarly to the asymmetric spin-boson case the agreement with the HEOM is excellent. The error for the entire 50 hold-out trajectories is $\epsilon=4.3\times 10^{-4}$. An even lower error in this case is expected as the dynamics of the symmetric spin-boson model is less rich compared to the asymmetric case and, thus, is easier to learn.
@@
In summary, we showed that a transformer model trained on a set of time-discretized values of the population difference of an open quantum system is capable of predicting the future time-evolution of a given trajectory with high accuracy across the non-Makorvian and strong coupling regimes. Only a short-time trajectory generated using a numerically accurate method is required, and the ML model can predict the long-time dynamics. This is the first implementation of the transformer model for the long-time dynamics of an open quantum system. It is yet another example of the ability of a ML model to predict complex physical phenomena with lower computational cost and high accuracy. The self-attention mechanisms are valuable tools that outperform many popular time-forecasting models such as RNNs. The value of the self-attention mechanics is in the ability to extract long-range correlations between the time points, in contrast to RNN which have limited memory. Additionally, the transformer layer can be executed in parallel enabling more efficient and faster training. Our transformer-based model is already accurate for the spin-boson Hamiltonian but its extensions to more complex (e.g., multi-level) systems might require using more elaborate architectures incorporating complex decoder, such as another transformer layer that is masked for future time points or a RNN such as long short-term memory (LSTM) or GRU.
@@
This work was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, under Award Number DE-SC0024511. A.A.K. also acknowledges NVIDIA Academic Hardware Grant Program.
@@
Codes used to generate results that are reported in the paper can be found in the following GitHub repository: https://github.com/kananenka-group/Transformer-spin-boson. The data set used in this work was published elsewhere and can be accessed at https://figshare.com/s/ed24594205ab87404238.

### v2

@@ Abstract
$\lambda/\Delta=\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0\},\;\omega_{c}/\Delta=\{1,2,3,4,5,6,7,8,9,10\}$, and $\beta\Delta=\{0.1,0.25,0.5,0.75,1\}$. HEOM method implemented in QuTiP software packageJohansson, Nation, and Nori (2012) was used in all calculations. The total propagation time was $t_{\max}\Delta=20$. The integration time-step was set to $t\Delta=0.05$. In total, 1000 HEOM calculations, 500 for symmetric $(\epsilon/\Delta=0)$ and 500 for asymmetric $(\epsilon/\Delta=1)$ spin-boson Hamiltonian were performed. Time-evolved subsystem reduced density matrices were saved every $\mathrm{d}t\Delta=0.1$. The population differences $\left\langle\hat{\sigma}_{z}(t)\right\rangle$ are calculated from the reduced denisty matrices and processed into shorter sequences of length $T$ by window slicing. Herrera Rodríguez and Kananenka (2021); Ullah and Dral (2021); Lin et al. (2021) Namely, for a time series $\mathbf{x}=\left(x^{(1)},\ldots,x^{(L)}\right)$, where $\left\langle\hat{\sigma}_{z}(t^{(j)})\right\rangle$ is denoted by $x^{(j)}$ for compactness, a slice is a subset of the original time series defined as $\mathbf{s}_{i:j}=\left(x^{(i)},\ldots,x^{(j)}\right)$ where $1\leqslant i\leqslant j\leqslant P$. For a given time series $\mathbf{x}$ of length $L$, and the length of the slice $P$, a set of $L-P+1$ sliced time series $\left\{\mathbf{s}_{1:P},\mathbf{s}_{2:P+1},\ldots,\mathbf{s}_{L-P+1:L}\right\}$ was generated. This procedure was also applied to the times $\mathbf{t}$ of the original sequence. Finally, the total data set $\mathcal{D}=\left\{\left(\{\mathbf{x}_{i},\mathbf{t}_{i}\},y_{i}\right)\right\}_{i=1}^{N}$ containing time series $\mathbf{x}_{i}$ and $\mathbf{t}_{i}$ with their corresponding labels $y_{i}$ is obtained by setting $1,\ldots,T$ elements of each slice, with $T=P-1$, to an input time-series $\mathbf{x}_{i}$ and the last ($P$th) element of each slice to the associated label $y_{i}$.
@@
In general, the size of the window $P-1$ or, equivalently $T$, should be treated as a hyperparameter but, following our previous works, Herrera Rodríguez and Kananenka (2021); Rodríguez et al. (2022) we set $T=0.2L$. Window slicing is applied to all 1000 HEOM reduced density matrices with different system and system-bath parameters. For each set of parameters, the initially calculated set of time-evolved $\left\langle\hat{\sigma}_{z}(t)\right\rangle$ with $L=t_{\max}/\mathrm{d}t=200$ generates 160 data points for the data set with $T=41$ (including $t\Delta=0$ point).
@@
From the raw HEOM data set of 1000 trajectories, 100 randomly chosen trajectories are taken as the hold-out test set, which is used for testing and generating the results. The remaining set of 900 trajectories is transformed into 144,000 trajectories by window slicing as described above. In total 72,000 short-time $\left\langle\hat{\sigma}_{z}(t)\right\rangle$ trajectories for symmetric and 72,000 for asymmetric spin-boson models were generated. Each trajectory has a time length of $t\Delta+\mathrm{d}t\Delta=4.1$ where $t\Delta=4.0$ is a part of a trajectory used as input vector $\mathbf{x}_{i}$. This length of an input trajectory was chosen to allow for a direct comparison to previously published ML models that use the same input.Herrera Rodríguez and Kananenka (2021); Rodríguez et al. (2022) As we showed previously, the longer the input trajectory the more accurate the prediction.Herrera Rodríguez and Kananenka (2021) However, generating a longer input trajectory is more computationally demanding. Therefore when choosing this parameter one has to balance the tradeoff between the cost of generating input data and the prediction accuracy. The last point $t\Delta+\mathrm{d}t\Delta=4.1$ is used as a label $y_{i}$. Following previous similar works the input data was not normalized.Herrera Rodríguez and Kananenka (2021); Ullah and Dral (2021) This data set of supervised trajectories is the training set which is randomly partitioned into two subsets: a sub-training set, which contains $80\%$ of the data and a validation set containing $20\%$ of the data. The transformer model parameters, weights and biases, were initially fitted on the sub-training set and the validation set is used for monitoring the performance of the models (mainly to prevent overfitting).
@@
Transformer Model. The transformer-based model employed here is inspired by the “Astromer” model.Donoso-Oliva et al. (2023) The model architecture is shown in Fig. 3. The input is the fixed-length trajectories corresponding to the population difference x and the corresponding time values t. In order to use the self-attention mechanism, the inputs are embedded in objects $X\in\mathbb{R}^{T\times d_{p}}$ that represent the original information of the time series in a $d_{p}$ dimensional space. After the self-attention block the input $X$ will be transformed into $Z\in\mathbb{R}^{T\times d_{p}}$. The population differences are projected into the $d_{p}=1\times 64$ space using a linear position-wise feed-forward neural network (PW-FFNN) without hidden layers. This operation transforms the population difference into a vector of size $d_{p}$. The time values are projected using the PE. At each step the PE is calculated using Eq. (8) projecting the time value into a vector of $d_{p}$. Note the match between the projected dimension of the FFNN for the population difference and the PE, doing so the two inputs are encoded in the space of the same dimensionality, and can be added as a single input for the self-attention layer. The FFNN learns how to encode the values of the population difference without interfering with the constant part of the PE. This means that the population difference will be represented in the higher dimension of $X=P+PE$ since the higher dimensions of the PE are constant as shown in Fig 2. After the addition, the inputs will be represented by the matrix $X$ of $41\times 64$.
@@
After the transformer layers, in order to reduce the dimensions before flattening, a PW-FNN is applied. After flattening, two fully connected FFNNs are used with a number of neurons of 1024 and 1408 respectively and RELU as the activation function. The final output of the ML model is a single neuron that gives the population difference at the next time step $\langle\hat{\sigma}(T+1)\rangle$.
@@
The model was implemented using Keras Chollet et al. (2015) software with TensorFlow blackened.Abadi et al. (2015) The model was trained using the adaptive moment estimation (Adam) algorithm Diederik (2014) with an initial learning rate of $1.0\times 10^{-4}$. The batch size was set to $N_{b}=128$ and the mean square error (MSE) was used as the loss function. The hyperparameters of the model: ##\##heads, $d_{p}$, the number transformer layers, and neurons in the fully-connected layers were optimized on the hold-out set for the asymmetric spin-boson model using KerasTunerO’Malley et al. (2019) for 50 iterations of Bayesian optimization. 500 epochs were used in the training. After the best model was chosen it was further trained for up to 4,500 epochs and the model with the lowest validation loss during the training was used. The MSE of the validation set was below $10^{-6}-10^{-7}$. This model was used to predict the population dynamics for the asymmetric spin-boson model.
@@
To generate the population dynamics for the symmetric spin-boson model, we retrained the model described above (for the asymmetric spin-boson model) based on the data set for the symmetric spin-boson Hamiltonian. The same hyperparameters and the initial weights were taken from the already trained asymmetric model. Since the asymmetric model already learned an adequate representation of input trajectories for the asymmetric spin-boson model, the symmetric ML model was trained only for 1000 epochs and the best model was chosen again based on validation MSE. The training was performed on a single NVIDIA RTX A6000 graphics card.
@@
Starting from the input sequence of $\mathbf{x}=\left(x^{(1)},\ldots,x^{(T)}\right)$ one uses a ML model to generate the population difference at $T+1$ (or $P$th) time step $y$. The newly predicted population difference is then combined with $T$ previous population differences to form another input vector $\mathbf{x}=\left(x^{(2)},\ldots,x^{(T)},y\to x^{(T+1)}\right)$ which is fed into the model again to generate the next prediction $y\to x^{(T+2)}$. Repeating this procedure will generate the long-time population dynamics from the initial short-time input. It should be noted that the next time value should match the time sequence, i.e. any time point should be $t^{(n)}=(n-1)dt$, and it is necessary to update the time input correspondingly.
@@
Figure 4: Expectation values $\left\langle\hat{\sigma}_{z}(t)\right\rangle$ for the asymmetric spin-boson model with $\epsilon=1$ as a function of time. Results predicted by the transformer-based ML model developed in this work (red circles) are compared to the numerically exact HEOM results (blue) for the following parameters: (a) $\lambda=0.1,\omega_{c}=6.0,\beta=0.75$; (b) $\lambda=0.3,\omega_{c}=8.0,\beta=1.0$; (c) $\lambda=0.2,\omega_{c}=10.0,\beta=0.25$; (d) $\lambda=0.4,\omega_{c}=8.0,\beta=0.75$; (e) $\lambda=0.8,\omega_{c}=10.0,\beta=1.0$; (f) $\lambda=0.7,\omega_{c}=10.0,\beta=0.1$. All parameters are in the units of $\Delta$.
In Fig. 4 we show the predictions of the population difference of our transformer model for the asymmetric spin-boson system for six representative sets of parameters. The agreement between ML-generated dynamics and numerically exact HEOM dynamics is excellent. The average error (MAE) between the predicted and reference trajectories for the entire hold-out set of 50 diverse reduced density matrix trajectories is $7.45\times 10^{-3}$. We stress that only a short trajectory of length $t\Delta=4$ is used as an input and these trajectories were not present in the training data set. To put this value into perspective, the error achieved in this work with the transformer architecture is lower than the lowest error of any NN-based models trained for the same system in our previous comprehensive study of ML models for long-time quantum dissipative dynamics.Rodríguez et al. (2022) There the lowest error of $2.14\times 10^{-2}$ was achieved for a convolutional gated recurrent unit (CGRU) model. In that work we also established that KRR models can be more accurate than their neural network counterparts. Importantly, the performance of our transformer-based model is now at the level of KRR models yet by using neural networks we can avoid the known issues associated with KRR models such as poor scaling with the amount of the training data. However, it is important to note that the number of trainable parameters in the case the transformer-based model presented here is $1,918,018$ which is four times the number of parameters of the convolutional GRU model. Increasing the number of parameters, however, does not necessarily guarantee the improvement in the performance of the model.
Figure 5: Expectation values $\left\langle\hat{\sigma}_{z}(t)\right\rangle$ for the symmetric spin-boson model as a function of time. Results predicted by the transformer-based ML model developed in this work (red circles) are compared to the numerically exact HEOM results (blue) for the following parameters: (a) $\lambda=0.2,\omega_{c}=8.0,\beta=1.0$; (b) $\lambda=0.4,\omega_{c}=10.0,\beta=1.0$; (c) $\lambda=0.2,\omega_{c}=10.0,\beta=0.25$; (d) $\lambda=0.1,\omega_{c}=4.0,\beta=0.1$; (e) $\lambda=0.8,\omega_{c}=3.0,\beta=1.0$; (f) $\lambda=1.0,\omega_{c}=2.0,\beta=0.1$. All parameters are in the units of $\Delta$.
@@
Fig. 5 illustrates the results for the symmetric spin-boson model. Similarly to the asymmetric spin-boson case the agreement with the HEOM is excellent. The error for the entire 50 hold-out trajectories is $4.3\times 10^{-4}$. An even lower error in this case is expected as the dynamics of the symmetric spin-boson model is less rich compared to the asymmetric case and, thus, is easier to learn.
@@
In summary, we showed that a transformer model trained on a set of time-discretized values of the population difference of an open quantum system is capable of predicting the future time-evolution of a given trajectory with high accuracy across the non-Makorvian and strong coupling regimes. Only a short-time trajectory generated using a numerically accurate method is required and our ML model can predict the long-time dynamics. This is the first implementation of the transformer model for the long-time dynamics of an open quantum system. It is yet another example of the ability of a ML model to predict complex physical phenomena with a lower computational cost compared to physics-based method and high accuracy. The self-attention mechanism is a valuable tool that outperforms many popular time-forecasting models such as RNNs. The value of the self-attention mechanics is in the ability to extract long-range correlations between the time points, in contrast to RNN which have limited memory. Additionally, the transformer layer can be executed in parallel, in contrast to the sequential of RNNs, enabling more efficient and faster training. Our transformer-based model is already accurate for the spin-boson Hamiltonian but its extensions to more complex (e.g., multi-level) systems might require using more elaborate architectures incorporating a complex decoder, such as another transformer layer that is masked for future time points or a RNN such as long short-term memory (LSTM) or GRU.
@@
This work was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under Award Number DE-SC0024511. L.E.H.R. thanks La Serena School for Data Science: Applied Tools for Astroinformatics, Biomedical Informatics, and Other Data-driven Sciences funded by NSF (AST-1637359). A.A.K. also acknowledges NVIDIA Academic Hardware Grant Program.
@@
The source code used in this study is available at https://github.com/kananenka-group/Transformer-spin-boson under release v1.0. The data set used in this work was published elsewhere and can be accessed at https://figshare.com/s/ed24594205ab87404238.